# CLIP-UP: A Simple and Efficient Mixture-of-Experts CLIP Training Recipe with Sparse Upcycling

### v1

@@ 4.1 Recipe Study
We first investigate the impact of using shared versus separated backbones and the effect of training from scratch compared to employing sparse upcycling. For a fair comparison, the shared backbone configuration uses 16 experts, matching the total number of experts (8 each for text and image) in the separated backbone setup. All experiments are conducted with the CLIP-B/16 model. All results of these models on ImageNet zero-shot classification and COCO/Flickr30k image-text retrieval tasks are illustrated in table 1.
The results show that the separated backbone combined with sparse upcycling delivers the best overall performance, achieving the highest ImageNet Top-1 accuracy of 76.9% , highest Flickr30K T2I and I2T recall@1 of 80.9% and 92.3%, and competitive results in COCO. Separated backbones generally outperform shared ones, likely because of the additional parameters dedicated to each modality. Notably, sparse upcycling provides greater improvements to shared backbones than to separated ones. Compared to shared models trained from scratch, sparse upcycling boosts ImageNet accuracy by 5.5% and COCO/Flickr retrieval by up to 7.1%, whereas the separated backbone achieves gains of 2.4% on ImageNet and up to 4.1% on retrieval tasks. Sparse upcycling proves effectiveness across both shared and separated architectures, consistently outperforming models trained from scratch. These results highlight the versatility and efficiency of CLIP-UP across different backbone configurations.
Figure 4: Performance comparison along with training EFLOPS of CLIP-UP and training from scratch on CLIP B/16.
We further explore the impact of adding LIMOE specific auxiliary loss described in Section 3. Specifically, we set $\tau=6$ to incentivize at least 6 experts (uniformly) or more (not necessarily in a uniform way) (Mustafa et al., 2022) are used for each modality and fine-tuned the loss weight under our experiment setup. As shown in Figure 3, applying the LIMOE auxiliary loss on the shared backbone trained from scratch significantly improves both ImageNet classification and COCO retrieval metrics, which is aligned with the finding from original LIMOE paper (Mustafa et al., 2022). However, its performance remains below other configurations without applying LIMOE specific auxiliary losses. Adding LIMOE specific loss to best configuration (Separated-Upcycle) achieves a slightly better text-image retrieval metrics but falls short on ImageNet zero-shot classficaition.
In addition, we also find adding these new auxiliary losses make the model hard to train as there are much more hyper-parameters to tune. To our best effort, we cannot make LIMOE specific auxiliary loss work in all configurations, e.g. in Shared-Upcycle and Separated settings. Due to the limit of resources, we decide to use the loss in equation 3.2 as the final training objective to reduce the effort of finding best hyperparameters.
@@ 4.2 Final Model Evaluation and Baselines
According to the results from the previous section, we select the separted backbone with sparse upcycling as the default recipe and further study its effectiness by training CLIP-UP models in different model sizes, from B/32 to L/14. Table 2 compares the zero-shot image classification and retrieval performance of CLIP-UP against baseline dense CLIP model trained for an equivalent number of steps across different scales. Extending training steps for dense CLIP from 440k to 790k yields only minor gains. In contrast, CLIP-UP demonstrates substantial improvements across all metrics. These gains are consistent across model scales, particularly for image-text retrieval tasks. Remarkably, CLIP-UP B/32, using only 47% of the inference GFLOPS, outperforms the baseline dense CLIP B/16 by 2.4% in COCO T2I recall@1 and matches its I2T recall@1. Similarly, CLIP-UP B/16, requiring just 31% of the inference GFLOPS, surpasses the baseline dense CLIP L/14 by 1.9% in COCO T2I recall@1 while matching its I2T recall@1. This highlights the effectiveness of training high-performance CLIP models by leveraging pretrained dense models through sparse upcycling with MoE.
@@ 4.3 Training Efficiency
To better illustrate the effectiveness of upcycle training, we also show the training process from dense pretraining + upcycling compared to training from scratch in Figure 4. The initial dense model training provides a strong foundation for subsequent upcycling. In contrast, training from scratch without leveraging the pretrained dense checkpoint, results in much higher computational costs to reach the performance level of CLIP-UP. This difference is especially noticeable for COCO image-to-text and ImageNet classification tasks. While sparse upcycling introduces an initial quality drop on ImageNet due to model reconfiguration (Komatsuzaki et al., 2023), CLIP-UP consistently outperforms the model trained from scratch, highlighting its efficiency in reducing training resources and improving overall performance.
5 Discussion and More Ablations
In this section, we discuss important architecture and training ablations relative to the configuration used during experiment. We use the CLIP B/16 with 8 experts as the default configuration here, expert capacity factor $C=2.0$, and 6 MoE layers interspersed in the Transformer for both modalities. The hyper-parameters are same as Section 4.
5.1 MoE added to single or multiple modalities
5.2 Expert capacity factor
Figure 6: Effects of different expert capacity $C$.
@@ 6 Conclusions
In this work, we introduce CLIP-UP, a simple yet efficient training strategy for CLIP that integrates MoE with sparse upcycling. Through extensive experiments on various configurations and auxiliary losses, we show that CLIP-UP significantly reduces training costs while consistently improving retrieval and classification performance across multiple model scales. Notably, CLIP-UP outperforms larger dense models in some tasks while drastically cutting inference FLOPs, highlighting sparse upcycling as a practical and scalable approach for building efficient, high-performance CLIP models. Additionally, ablation studies provide comprehensive insights into key design decisions.
Acknowledgement
@@ A.1 Training hyper-parameters
We primarily follow (Radford et al., 2021) for hyper-parameter selection, using the WIT-3000M (Wu et al., 2024) and DFN-5B (Fang et al., 2023) training datasets. Table 3 summarizes the hyper-parameters for all experiments, including MoE-specific configurations and parameters for dense CLIP, sparse CLIP, and CLIP-UP.
@@
Text tokenizer T5 (Raffel et al., 2023), lowercase
@@
Inner structure Pre-Layer Normalization (Xiong et al., 2020)
@@ Appendix B Ablation study
B.1 Normalize gating weights before or after routing.
@@ C.1 Comparison of model architectures and impact of LIMOE auxiliary loss
C.2 Comparison of MoE added to single modality or both modalities
@@
C.3 Comparison of capacity factor
All results from Figure 6 to compare the impact of capacity factor $C$.
@@ C.3 Comparison of capacity factor
Model Imagenet COCO Flickr30K
Accuracy@1 T2I R@1 I2T R@1 T2I R@1 I2T R@1
$C_{image}=2,C_{text}=2$ 76.9 52.1 71.5 80.9 92.3
$C_{image}=4,C_{text}=2$ 78.4 46.3 66.9 75.5 90.2

### v2

@@ 3.1 Recipe Study
We compare shared vs. separated backbones and training from scratch vs. sparse upcycling using the CLIP-B/16 model, with the shared setup using 16 experts to match the separated configuration (8 per modality). As shown in Table 1, the separated backbone with sparse upcycling delivers the best overall performance due to dedicated parameters per modality, while the shared backbone sees greater relative gains from sparse upcycling. Overall, sparse upcycling consistently outperforms training from scratch, demonstrating CLIP-UP’s versatility and efficiency across configurations.
3.2 Impact of LIMOE auxiliary loss.
Figure 3: Impact of LIMOE auxiliary loss under different training setups. Adding LIMOE loss sometimes causes instability, especially with unshared backbones, while our upcycling recipe remains more robust.
We examine the LIMOE auxiliary loss by setting $\tau=6$ in the global entropy loss Mustafa et al. (2022), encouraging use of at least six experts per modality, and tuning the loss weight for our setup. As shown in Figure 3, it improves ImageNet and COCO performance with a shared backbone trained from scratch, consistent with prior work  Mustafa et al. (2022), but still underperforms compared to other configurations without it. Applying the loss to our best setup (Separated-Upcycle) slightly boosts text-image retrieval but falls short on ImageNet zero-shot classification.
These auxiliary losses also increase training complexity due to more hyperparameters. As LIMOE loss didn’t work reliably across all setups, we use load balance and router-$z$ losses to simplify tuning under resource constraints.
@@ 3.3 Final Model Evaluation and Baselines
Based on previous results, we adopt the separated backbone with sparse upcycling as the default setup and evaluate CLIP-UP across model sizes from B/32 to L/14. Table 2 compares zero-shot classification and retrieval performance against dense CLIP models trained for the same number of steps. While extending dense CLIP training from 440k to 790k steps yields minor gains, CLIP-UP shows consistent, significant improvements across scales, especially in retrieval. Notably, CLIP-UP B/32 uses only 47% of the inference GFLOPS yet outperforms dense CLIP B/16 in COCO T2I recall@1 by 2.4%, while CLIP-UP B/16 uses just 31% of the GFLOPS and surpasses dense CLIP L/14 by 1.9%. These results demonstrate the efficiency and effectiveness of sparse upcycling for scaling CLIP models.
Figure 4: Performance vs. training EFLOPS for CLIP-UP and sparse-from-scratch model on CLIP B/16.
@@ 3.4 Training Efficiency
To highlight the effectiveness of upcycle training, Figure 4 compares dense pretraining + upcycling with training from scratch. The pretrained dense model provides a strong starting point, while training from scratch requires significantly more compute to match CLIP-UP’s performance—especially on COCO image-to-text and ImageNet. Although sparse upcycling initially causes a performance drop on ImageNet due to reconfiguration, CLIP-UP consistently outperforms the scratch baseline, demonstrating better efficiency and overall performance.
@@ 4 Conclusions
We present CLIP-UP, an efficient CLIP training strategy that combines MoE with sparse upcycling. Extensive experiments show it reduces training costs and inference FLOPs while improving performance across scales, even outperforming larger dense models. Ablation studies shown in Appendix B further validate key design choices, highlighting CLIP-UP’s practicality and scalability.
5 Limitation
While our proposed method demonstrates strong performance improvements in retrieval tasks such as COCO and Flickr30K, it reveals a trade-off with classification performance on ImageNet and its variants. Specifically, we have not yet identified a training configuration that yields significant gains across both retrieval and classification simultaneously. Our current best setup prioritizes retrieval effectiveness, achieving notable improvements on COCO and Flickr30K, but leads to only marginal gains on ImageNet.
We discuss this trade-off in more detail in Appendix B.3, highlighting the role of the expert capacity factor in shaping task-specific performance. In particular, we provide examples showing how ImageNet and COCO respond differently to token dropping under varying expert capacities, which we believe contributes to the observed trade-off. While these insights offer a preliminary understanding, we are still exploring more effective strategies to better balance retrieval and classification performance.
6 Acknowledgement
@@ A.1 Training hyper-parameters
We primarily follow Radford et al. (2021) for hyper-parameter selection, using the WIT-3000M Wu et al. (2024) and DFN-5B Fang et al. (2023) training datasets. Table 3 summarizes the hyper-parameters for all experiments, including MoE-specific configurations and parameters for dense CLIP, sparse CLIP, and CLIP-UP.
@@
Text tokenizer T5 Raffel et al. (2023), lowercase
@@
Inner structure Pre-Layer Normalization Xiong et al. (2020)
@@ Appendix B Ablation study
B.1 MoE added to single or multiple modalities
B.2 Expert capacity factor
Figure 6: Effects of different expert capacity $C$.
B.3 Normalize gating weights before or after routing.
@@ C.1 Comparison of model architectures and impact of LIMOE auxiliary loss
C.2 Performance on ImageNet variants
To complement the performance comparison in Table 2, we additionally evaluated our models on several ImageNet variants, including ImageNet-V2 Recht et al. (2019), ImageNet-A Hendrycks et al. (2021b), and ImageNet-R Hendrycks et al. (2021a). As shown in Table 5, the performance trends on these datasets are consistent with those observed on the original ImageNet benchmark.
Table 5: Performance on ImageNet Variants of CLIP-UP, dense models across model sizes.
Model Steps Imagenet-V2 Imagenet-A Imagenet-R
classification classification classification
(k) Acc@1 Acc@5 Acc@1 Acc@5 Acc@1 Acc@5
B/32
CLIP (ours) 440 64.0 88.0 32.1 65.4 80.0 92.7
CLIP (ours) 790 64.0 87.7 32.6 64.8 80.3 92.7
CLIP-UP 790 65.1 88.5 34.2 66.8 79.9 92.8
B/16
CLIP (ours) 440 68.2 90.7 47.7 78.5 84.6 95.3
CLIP (ours) 790 69.3 91.2 50.5 80.2 85.9 95.8
CLIP-UP 790 69.6 91.1 49.2 78.6 85.9 95.6
L/14
CLIP (ours) 440 74.3 93.5 67.1 89.0 90.6 97.7
CLIP (ours) 790 74.7 93.6 68.2 88.9 91.2 97.8
CLIP-UP 790 73.8 93.6 66.1 88.1 89.7 97.3
C.3 Comparison of MoE added to single modality or both modalities
@@
C.4 Comparison of capacity factor
As discussed in Section 6, expert capacity factor $C$ plays a crucial role in balancing classification performance on ImageNet and its variants with retrieval performance on Flickr and COCO. As shown in Table 7, increasing $C$ from 2.0 to 4.0 consistently improves accuracy on ImageNet and its variants, but at the expense of COCO performance—highlighting an inherent trade-off.
@@ C.4 Comparison of capacity factor
Model Imagenet Imagenet-V2 Imagenet-A Imagenet-R COCO Flickr30K
Acc@1 Acc@1 Acc@1 Acc@1 T2I R@1 I2T R@1 T2I R@1 I2T R@1
$C_{image}=2,C_{text}=2$ 76.9 69.4 49.8 83.9 52.1 71.5 80.9 92.3
$C_{image}=4,C_{text}=2$ 78.4 71.0 53.7 86.9 46.3 66.9 75.5 90.2
Appendix D Router Analysis
D.1 Routing distribution
The routing is balanced across all transformer layers. The average token ratios assigned to each expert for both text and image modalities on ImageNet and COCO are shown in Figure 9 and Figure 10, respectively.
Token assignment appears more balanced for images than for text, likely due to the higher number of tokens in images, many of which carry similar or redundant information. This redundancy facilitates a more uniform distribution of tokens across experts. In contrast, text is typically more discrete and information-dense, leading to a stronger preference for certain experts that specialize in specific linguistic patterns.
Figure 9: Visualization of the token assignment ratios to each expert in each layer by the router on ImageNet — image tokens (top) and text tokens (bottom).
Figure 10: Visualization of the token assignment ratios to each expert in each layer by the router on COCO — image tokens (top) and text tokens (bottom).
D.2 Expert preference pattern
We observe distinct preference patterns among experts. As illustrated in Figure 11, one expert predominantly processes tokens related to "eyes," consistently attending to the eye regions of various animals and humans. Another expert demonstrates specialization in "symbols," handling text tokens from a wide range of contexts, including posters, machine interfaces, book titles, store signage, and instructional materials.
Figure 11: Visualization of expert preference pattern examples. Red bounding boxes highlight the image tokens assigned to the expert.